# On the Scalability of GNNs for Molecular Graphs

### v1

@@ Abstract
Scaling deep learning models has been at the heart of recent revolutions in language modelling and image generation. Practitioners have observed a strong relationship between model size, dataset size, and performance. However, structure-based architectures such as Graph Neural Networks (GNNs) are yet to show the benefits of scale mainly due to the lower efficiency of sparse operations, large data requirements, and lack of clarity about the effectiveness of various architectures. We address this drawback of GNNs by studying their scaling behavior. Specifically, we analyze message-passing networks, graph Transformers, and hybrid architectures on the largest public collection of 2D molecular graphs. For the first time, we observe that GNNs benefit tremendously from the increasing scale of depth, width, number of molecules, number of labels, and the diversity in the pretraining datasets, resulting in a 30.25% improvement when scaling to 1 billion parameters and 28.98% improvement when increasing size of dataset to eightfold. We further demonstrate strong finetuning scaling behavior on 38 tasks, outclassing previous large models. We hope that our work paves the way for an era where foundational GNNs drive pharmaceutical drug discovery.
@@ 1 Introduction
As summarized in Figure 1, we aim to answer the above question by studying the scaling behavior of 2D molecular GNNs under different settings of width, depth, number of molecules, number of labels, and the diversity in datasets. We analyze message-passing networks, graph Transformers, and hybrid architectures on the largest public collection of 2D molecular graphs. All models are tested in 2 different settings: (1) randomly split train and test sets for pretraining and (2) fine-tuning of pretrained models on standard benchmarks.
@@
  • Finally, we show that, simply by scaling the model width upto 1B parameters, our largest model outperforms other pretrained models, while consistently achieving parity with the state-of-the-art specialized models across a wide variety of competitive downstream tasks.
@@ 5.2 Scaling Trends on Downstream Tasks
Figure 4: Comparison of our 1B Ensemble probing approach (that combines fingerprints from the MPNN++, Transformer and hybrid GPS++ model) to the SOTA across TDC, Polaris, and MoleculeNet benchmarks. For TDC, a group of 8 different models establishes SOTA across the 22 tasks. Ensemble probing establishes new SOTA on $12/22$ TDC tasks and on all but two tasks among Polaris and MoleculeNet.
@@
5.3 Comparison to Other Models
We now present two further optimized probing approaches that enhance performance on the various downstream tasks considered above, while additionally considering the MoleculeNet benchmark (Wu et al., 2018) here. For each benchmark in MoleculNet, we compare to SOTA performance as of March 15, 2024 and to the MolE model (Méndez-Lucio et al., 2022) if results are available. MolE is a particularly relevant baseline as, similar to our approach, it can be flexibly applied to many downstream tasks without major modifications, while SOTA across the benchmarks is naturally established by various different approaches depending on the specific task.
@@ 5.3 Comparison to Other Models
Figure 5: Comparison of our GPS++ probing (that leverages multiple fingerprints) and the 1B Ensemble probing (that leverages fingerprints from MPNN++, Transformer, and the hybrid GPS++) to various baselines across TDC benchmark collection using an aggregated metric (average standardized mean of the leaderboard).
Ensemble probing. We also found that downstream task performance can be further enhanced by ensembling fingerprints from multiple pretrained models. Our ensemble probing approach extracts fingerprints from 1B MPNN++, Transformer, and GPS++ models, followed by an MLP. Figure 4 compares this model across the TDC, Polaris and MoleculeNet benchmark collections to the current SOTA for each task and to MolE. Our ensemble probing approach yields the strongest downstream task results, outperforming MolE in $19/22$ TDC tasks and establishing SOTA performance on $12/22$ TDC tasks. This makes our ensemble probing approach the model with the most SOTA entries in the TDC leaderboard followed by MapLight+GNN (Notwell & Wood, 2023) that established SOTA on 5 TDC tasks and 7 other methods that are SOTA for at least one TDC task. [^3] We report scores from the TDC leaderboards extracted on March 15, 2024. SOTA on TDC is established by a group of 8 different models, namely Chemprop-RDKit (Yang et al., 2019), MapLight, MapLight+GNN (Notwell & Wood, 2023), BaseBoosting (Huang12 et al., ), CFA (Quazi et al., 2023), SimGCN, ZairaChem (Oselusi et al., 2024) and ContextPred (Zhou et al., 2020). Finally, when comparing the ensemble probing to previous best methods on the Polaris and MoleculeNet benchmarks, we observe that our model is significantly better (often by large margins). We primarily attribute the large-scale success of our model to width scaling up to 1B parameters, suggesting us to scale even further.
We also report the aggregated score of the 1B ensemble probing in Figure 5, which almost reaches the aggregated SOTA performance, which is remarkable recalling again that this score is a derived from the best scoring method per task of the benchmark collection, while we use a single method for all tasks.
@@ 6 Conclusion
In this paper, we studied the scalability of GNN models including message-passing networks, graph Transformers and hybrid architectures on the largest public collection of 2D molecules for the tasks of molecular property prediction. We showed major performance gains from the growing amount of parameters, data and compute, both on the original test set and on downstream finetuning. Importantly, our models benefit tremendously from the increasing scale of width, number of molecules, and number of labels. Our largest 1B parameter models, including MPNN++, Transformer, and GPS++, continue to scale favourably resulting in peak 60% improvement across precision and correlation metrics. More importantly, we demonstrate a consistent performance improvement on downstream property prediction tasks. Finetuned 1B parameter models consistently produce results on par with specialized state-of-the-arts. We hope that our work paves the way for the development of foundational GNNs and new architectures with applications in pharmaceutical advancements and drug discovery.

### v2

@@ Abstract
Scaling deep learning models has been at the heart of recent revolutions in language modelling and image generation. Practitioners have observed a strong relationship between model size, dataset size, and performance. However, structure-based architectures such as Graph Neural Networks (GNNs) are yet to show the benefits of scale mainly due to the lower efficiency of sparse operations, large data requirements, and lack of clarity about the effectiveness of various architectures. We address this drawback of GNNs by studying their scaling behavior. Specifically, we analyze message-passing networks, graph Transformers, and hybrid architectures on the largest public collection of 2D molecular graphs. For the first time, we observe that GNNs benefit tremendously from the increasing scale of depth, width, number of molecules, number of labels, and the diversity in the pretraining datasets. We further demonstrate strong finetuning scaling behavior on 38 highly competitive downstream tasks, outclassing previous large models. This gives rise to MolGPS, a new graph foundation model that allows to navigate the chemical space, outperforming the previous state-of-the-arts on 26 out the 38 downstream tasks. We hope that our work paves the way for an era where foundational GNNs drive pharmaceutical drug discovery.
@@ 1 Introduction
As summarized in Figure 1, we aim to answer the above question by studying the scaling behavior of 2D molecular GNNs under different settings of width, depth, number of molecules, number of labels, and the diversity in datasets. We analyze message-passing networks, graph Transformers, and hybrid architectures on the largest public collection of 2D molecular graphs. The models are tested in 2 different settings: (1) randomly split train and test sets for pretraining and (2) finetuning/probing of pretrained models on standard benchmarks.
@@
  • Finally, we propose MolGPS, a foundation model derived from our findings on how to best scale molecular GNNs. MolGPS constitutes the most dominant model across the presented benchmarks to date, establishing state-of-the-art (SOTA) on 26/38 highly competitive downstream tasks.
@@ 5.2 Scaling Trends on Downstream Tasks
Figure 4: Comparison of our MolGPS foundation model (that combines fingerprints from the MPNN++, Transformer and hybrid GPS++ model) to the SOTA across TDC, Polaris, and MoleculeNet benchmarks. For TDC, a group of 8 different models establishes SOTA across the 22 tasks. MolGPS establishes new SOTA on $12/22$ TDC tasks and on all but two tasks among Polaris and MoleculeNet.
@@
5.3 Towards a final foundation model
We now explain how the above findings can be pieced together to develop a highly competitive graph foundation model. Apart from scaling the model width, we found two other design choices that had a major impact on the performance for the various downstream tasks. We additionally report results on the MoleculeNet benchmark (Wu et al., 2018) here in addition to the previously used TDC and Polaris benchmarks.
Multi-fingerprint probing. Our previous task-head ablation study suggested that different fingerprints may be optimal for probing depending on the downstream task. As a result we further experimented with probing on combinations of multiple fingerprints extracted at different layers of a pretrained model, which improves performance on downstream tasks. Results can be further enhanced by combining fingerprints from multiple pretrained models.
Pretraining without L1000. Additionally, based on our observation in the dataset ablation, we pretrained new models without the L1000_* pretraining tasks, which leads to performance improvements across all scales. We hypothesize this is due to the challenging signal-to-noise ratio for those particular tasks, as also pointed out in the literature (Tong et al., 2023).
@@ 5.3 Towards a final foundation model
Figure 5: Comparison of our GPS++ probing (that leverages multiple fingerprints) and our MolGPS foundation model (that leverages fingerprints from MPNN++, Transformer, and the hybrid GPS++) to various baselines across TDC benchmark collection using an aggregated metric (average standardized mean of the leaderboard).
MolGPS. We now introduce a final graph foundation model that leverages the various findings of this paper. MolGPS inherits many architecture design choices from the General, Powerful, Scalable Graph Transformer method (Rampášek et al., 2022) and can be used to navigate the molecular space. MolGPS combines fingerprints from 1B MPNN++, Transformer, and GPS++ models that have been pretrained without the L1000 dataset, followed by an MLP. Figure 4 compares this model across the TDC, Polaris and MoleculeNet benchmark collections to the current SOTA for each task and to MolE. MolGPS yields by far the strongest downstream task results, outperforming MolE in $19/22$ TDC tasks and establishing SOTA performance on $12/22$ TDC tasks. This makes MolGPS the model with the most SOTA entries in the TDC leaderboard followed by MapLight+GNN (Notwell & Wood, 2023) that established SOTA on 5 TDC tasks and 7 other methods that are SOTA for at least one TDC task. [^3] We report scores from the TDC leaderboards extracted on March 15, 2024. SOTA on TDC is established by a group of 8 different models, namely Chemprop-RDKit (Yang et al., 2019), MapLight, MapLight+GNN (Notwell & Wood, 2023), BaseBoosting (Huang12 et al., ), CFA (Quazi et al., 2023), SimGCN, ZairaChem (Oselusi et al., 2024) and ContextPred (Zhou et al., 2020). Similarly, when compared to previous best methods on the Polaris and MoleculeNet benchmarks, we observe that our model is significantly better (often by large margins) for all but two downstream tasks. We primarily attribute the large-scale success of our model to width scaling up to 1B parameters, suggesting us to scale even further.
In Figure 5, we also report the aggregated score on the TDC benchmark for MolGPS, which almost reaches the aggregated SOTA performance, which is remarkable recalling again that this score is a derived from the best scoring method per task of the benchmark collection, while we use a single method for all tasks.
@@ 6 Conclusion
In this paper, we studied the scalability of GNN models including message-passing networks, graph Transformers and hybrid architectures on the largest public collection of 2D molecules for the tasks of molecular property prediction. We showed major performance gains from the growing amount of parameters, data and compute, both on the original test set and on downstream finetuning. Importantly, our models benefit tremendously from the increasing scale of width, number of molecules, and number of labels. Our largest 1B parameter models, including MPNN++, Transformer, and GPS++, continue to scale favourably. More importantly, we demonstrate a consistent performance improvement on downstream property prediction tasks via finetuning and probing. Finally, we derive MolGPS, a powerful foundational model based on a multi-fingerprint probing approach that can be used to navigate the chemical space, establishing state-of-the-art on 26 out of 38 highly competitive downstream tasks. We hope that our work paves the way for the development of foundational GNNs and new architectures with applications in pharmaceutical advancements and drug discovery.